OpenStudy (anonymous):

Two cars of masses m1 and m2 are moving in a circle of radii r1 and r2 , respectively. Their speeds are such that they make complete circles in the same time t.The ratio of their centripetal acceleration?

OpenStudy (anonymous):

@ajprincess

OpenStudy (shubhamsrg):

is the ans p1/p2 ? where p=momentum ?

OpenStudy (anonymous):

ac = v^2/r

OpenStudy (anonymous):

no the answer is m1 r1 : m2 r2

OpenStudy (anonymous):

i wonder hw that mass came there

OpenStudy (anonymous):

becoz centripetal force = mv^2/r.......but the question asks for centripetal acceleration

OpenStudy (shubhamsrg):

shouldnt ans be m1v1 / m2v2 ?

OpenStudy (shubhamsrg):

instead of r1 ?

OpenStudy (shubhamsrg):

and r2 *

OpenStudy (anonymous):

i am confused

OpenStudy (shubhamsrg):

see a1 = m1(v1)^2 / r1 and a2 = m2(v2)^2 /r2 you have to find a1/a2 = (m1v1 / m2v2) * ( v1r2 / v2r1) now t = 2pi r1/v1 = 2pi r2/v2 =>v1r2 = r1v2 so req ratio = m1v1 / m2v2

OpenStudy (anonymous):

first of all @shubhamsrg centripetal acceleration = v^2/r

OpenStudy (shubhamsrg):

ohh yes..sorry..

OpenStudy (shubhamsrg):

really sorry..

OpenStudy (shubhamsrg):

well v1 and v2 are not mentioned in the ques..and ans comes out to be v1/v2 we need to fine a relation between m,v and r now..hmm..

OpenStudy (shubhamsrg):

i dont see mass coming from any where.. since v1/v2 = r1/r2 i'd say thats the ans..

OpenStudy (anonymous):

Accelerations are independent of mass.It might be given to confuse you. Anyway,Since the cars take the same time to go around the circle, angular velocity = 2*pi/t = same for both cars. Let the centripetal accelerations be a1 and a2.v= r *angular velocity. a1 = v1^2/r1 = r1*(angular velocity)^2, a2 = v2^2/r2 = r2*(angular velocity)^2 a1/a2 = r1/r2 .

OpenStudy (anonymous):

@kartiksriramk but in the book it is given as m1 r1 : m2 r2

OpenStudy (anonymous):

Please check it again.I am pretty sure that the ratio is r1/r2.

OpenStudy (anonymous):

i also got that...... the options ar r1 : r2 , m1r1:m2r2

OpenStudy (anonymous):

Then it is definitely r1 : r2 .

OpenStudy (anonymous):

but the answer is given as m1r1:m2r2

OpenStudy (anonymous):

It must be wrong.

OpenStudy (anonymous):

i dont think

OpenStudy (anonymous):

i will show the proof

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

AIEEE question

OpenStudy (anonymous):

and this is a model answer paper

OpenStudy (anonymous):

Did you check the answer key?

OpenStudy (anonymous):

he may be wrong @kartiksriramk

OpenStudy (experimentx):

they have same angular velocity, hence the required result follows